OpenStudy (anonymous):

If you multiply x by x^2 do you get x^3?

OpenStudy (amistre64):

lets see: x^2 means: x*x x*x*x means x^3 .. so id say yes

OpenStudy (anonymous):

YUP.

OpenStudy (waheguru):

sorry chess machine plz i still need help i dod ont compleatly understand

OpenStudy (anonymous):

\[x^{a} \times x^{b}= x^{a+b}\]

OpenStudy (anonymous):

Thus, x^1 times X^2 = x^ (2+1)=x^3